# Large-scale distributed synchronization systems, using a cancel-on-completion redundancy mechanism

### v1

@@ 8 Proofs
Now, suppose the theorem statement does not hold. Then, there exists a sequence $x^{(\kappa)}$ of ML-FP, indexed by $\kappa\uparrow\infty$, along which $\Phi_{1}(x^{(\kappa)})\uparrow\infty$, and all other system parameters converge (to a set of parameters within the specified class). In this case, obviously, both the left boundary (if any) and right boundary (if any) must go to, respectively, $\infty$- ∞ or +$\infty$+ ∞. We now show that there exists $\epsilon>0$, such that for all sufficiently large $\kappa$,
@@
(Compare to (32).) To prove (41) we use the same argument as in the proof of (32) in Theorem 6.1, except we do not need Assumption 3.4 to claim the analog of property (38) in the case (a) (see text following (38)). This is because $x^{(\kappa)}_{w}$ are uniformly Lipschitz and, therefore, for a sufficiently small $\delta_{3}>0$, the $x^{(\kappa)}$-measure of the interval $[a^{+}_{\nu}-\delta_{3},a^{+}_{\nu}]$ is uniformly small.
@@
Now, let us consider a sequence of $w^{*}\downarrow-\infty$, and for each $w^{*}$ we consider the ML-FP $x$ in interval $(-\infty,w^{*}]$, which we shift to the right so that the right boundary is at 00, and normalize so that the scaled number of servers is $1$ instead of $1-x_{w^{*}}$. Namely, for each $w^{*}$, consider
@@
For any $v$ satisfying (24), and any $\epsilon\in(0,1)$ sufficiently close to $1$, the ML-MFP $x^{*}$ in Lemma 8.2 is the unique ML-FP for the $(\epsilon,0)$-reduced system with the right boundary at 00.
@@
The rest of the proof is by contradiction. Suppose that in addition to the ML-MFP $x^{*,\epsilon}$ for speed $v$, there exists a different ML-FP $x$ for the same speed. ML-FP $x$ is dominated by, but not equal to, $x^{*,\epsilon}$; also, by Lemma 8.2, $x$ has an asymptotically exponential left tail with the same exponent $\beta$ as $x^{*,\epsilon}$ does. Let us choose speed $\hat{v}>v$ sufficiently close to $v$, so that ML-MFP $\hat{x}^{*,\epsilon}$ for speed $\hat{v}$ is such that in some interval $\hat{x}^{*,\epsilon}_{w}>x_{w}$. The left tail of $\hat{x}^{*,\epsilon}$ is asymptotically exponential with exponent $\hat{\beta}<\beta$; this, in particular means that for all $w$ sufficiently far on the left (i.e. negative, with sufficiently large $|w|$) $\hat{x}^{*,\epsilon}_{w}<x_{w}$. Therefore, as $w$ increases from $\infty$- ∞ to 00, the difference $\hat{x}^{*,\epsilon}_{w}-x_{w}$ is first negative, then there is a point when it is strictly positive, and at the right boundary 00 it is 00. This situation is impossible by Lemma 5.14. $\Box$
Lemma 8.3 says that, for any $v$ satisfying (24), for any $\epsilon\in(0,1)$ sufficiently close to $1$, ML-FP for the $(\epsilon,0)$-reduced system with the right boundary at 00 is unique. But, by Lemma 5.11, any DE-FP $x$ is uniquely determined by its left tail, taken as initial condition. Therefore, as a corollary, we obtain the following
@@
Conclusion of the proof of Theorem 7.3. We see that, for any $v$ satisfying (24), there is a unique function $x^{*}$ which determines all ML-FP with proper left tail. Namely, choose $\epsilon^{\prime}$ sufficiently close to $1$, and consider the corresponding unique DE-FP $x^{*,\epsilon^{\prime}}$ in the interval $(-\infty,0]$, such that $x^{*,\epsilon^{\prime}}_{0}=\epsilon^{\prime}$. (This DE-FP $x^{*,\epsilon^{\prime}}$ is the rescaled version of the unique ML-FP = ML-MFP for the $(\epsilon^{\prime},0)$-reduced system with frame $(-\infty,0]$.) Continue this DE-FP solution to the positive half-axis. The solution will either (a) hit the $w$-axis or (b) will converge to some $\epsilon_{*}\geq 0$ as $w\uparrow\infty$. In case (a): for any $10\leq\epsilon<$ 1 there is a unique proper ML-FP for the $(\epsilon,0)$-reduced right-regulated system. In case (b): for any $\epsilon_{*}<\epsilon<1$ there is a unique proper ML-FP for the $(\epsilon,0)$-reduced right-regulated system; for $\epsilon=\epsilon_{*}$ there is a unique proper ML-FP the $(\epsilon,0)$-reduced free system; for $\epsilon<\epsilon_{*}$ there is a unique ML-FP with proper left tail, but not proper on the right, for the $(\epsilon,0)$-reduced free system. $\Box$
@@ 9 Discussion
In this paper we do identify one case when “all possible nice properties hold,” namely the case when Condition 6.5 holds for any speed $v$. (In turn, Condition 6.5 necessarily holds when all job classes are $D$-monotone. And, in turn, if all classes have i.i.d. component sizes with IHR, this is sufficient for $D$-monotonicity.) Our results show that in this “nice” case: (a) $v_{min}=v_{max}=v_{*}$ and, therefore, the free ML-FP is unique and has speed $v_{*}$; (b) SSAI for the left-regulated system holds for the full range of loads; (c) the average advance velocity $v_{n}$ of the free system (with speed parameter 00) converges to $v_{*}$. Our results do not rule out that Condition 6.5 for any speed $v$ always holds for our general model, and then properties (a)-(c) always hold. Or, it is possible that Condition 6.5 does not always hold, but the nice properties (a)-(c) still hold always. Addressing these questions may be a subject of future work.

### v2

@@ 8 Proofs
Now, suppose the theorem statement does not hold. Then, there exists a sequence $x^{(\kappa)}$ of ML-FP, indexed by $\kappa\uparrow\infty$, along which $\Phi_{1}(x^{(\kappa)})\uparrow\infty$, and all other system parameters converge (to a set of parameters within the specified class). In this case, obviously, both the left boundary (if any) and right boundary (if any) must go to, respectively, -$\infty$ or $\infty$. We now show that there exists $\epsilon>0$, such that for all sufficiently large $\kappa$,
@@
(Compare to (32).) To prove (41) we use the same argument as in the proof of (32) in Theorem 6.1, except we do not need Assumption 3.4 to claim the analog of property (38) in the case (a) (see text following (38)). This is because $x^{(\kappa)}_{w}$ are uniformly Lipschitz and, therefore, for a sufficiently small $\delta_{3}>0$, the $x^{(\kappa)}$-measure of the interval $a^{+}_{\nu}-\delta_{3},a^{+}_{\nu}$] is uniformly small.
@@
Now, let us consider a sequence of $w^{*}\downarrow$-$\infty$, and for each $w^{*}$ we consider the ML-FP $x$ in interval $(-\infty,w^{*}]$, which we shift to the right so that the right boundary is at 0, and normalize so that the scaled number of servers is $1$ instead of $1-x_{w^{*}}$. Namely, for each $w^{*}$, consider
@@
For any $v$ satisfying (24), and any $\epsilon\in(0,1)$ sufficiently close to $1$, the ML-MFP $x^{*}$ in Lemma 8.2 is the unique ML-FP for the $(\epsilon,0)$-reduced system with the right boundary at 0.
@@
The rest of the proof is by contradiction. Suppose that in addition to the ML-MFP $x^{*,\epsilon}$ for speed $v$, there exists a different ML-FP $x$ for the same speed. ML-FP $x$ is dominated by, but not equal to, $x^{*,\epsilon}$; also, by Lemma 8.2, $x$ has an asymptotically exponential left tail with the same exponent $\beta$ as $x^{*,\epsilon}$ does. Let us choose speed $\hat{v}>v$ sufficiently close to $v$, so that ML-MFP $\hat{x}^{*,\epsilon}$ for speed $\hat{v}$ is such that in some interval $\hat{x}^{*,\epsilon}_{w}>x_{w}$. The left tail of $\hat{x}^{*,\epsilon}$ is asymptotically exponential with exponent $\hat{\beta}<\beta$; this, in particular means that for all $w$ sufficiently far on the left (i.e. negative, with sufficiently large $|w|$) $\hat{x}^{*,\epsilon}_{w}<x_{w}$. Therefore, as $w$ increases from -$\infty$ to 0, the difference $\hat{x}^{*,\epsilon}_{w}-x_{w}$ is first negative, then there is a point when it is strictly positive, and at the right boundary 0 it is 0. This situation is impossible by Lemma 5.14. $\Box$
Lemma 8.3 says that, for any $v$ satisfying (24), for any $\epsilon\in(0,1)$ sufficiently close to $1$, ML-FP for the $(\epsilon,0)$-reduced system with the right boundary at 0 is unique. But, by Lemma 5.11, any DE-FP $x$ is uniquely determined by its left tail, taken as initial condition. Therefore, as a corollary, we obtain the following
@@
Conclusion of the proof of Theorem 7.3. We see that, for any $v$ satisfying (24), there is a unique function $x^{*}$ which determines all ML-FP with proper left tail. Namely, choose $\epsilon^{\prime}$ sufficiently close to $1$, and consider the corresponding unique DE-FP $x^{*,\epsilon^{\prime}}$ in the interval $(-\infty,0]$, such that $x^{*,\epsilon^{\prime}}_{0}=\epsilon^{\prime}$. (This DE-FP $x^{*,\epsilon^{\prime}}$ is the rescaled version of the unique ML-FP = ML-MFP for the $(\epsilon^{\prime},0)$-reduced system with frame $(-\infty,0]$.) Continue this DE-FP solution to the positive half-axis. The solution will either (a) hit the $w$-axis or (b) will converge to some $\epsilon_{*}\geq 0$ as $w\uparrow\infty$. In case (a): for any $0\leq\epsilon<1$ there is a unique proper ML-FP for the $(\epsilon,0)$-reduced right-regulated system. In case (b): for any $\epsilon_{*}<\epsilon<1$ there is a unique proper ML-FP for the $(\epsilon,0)$-reduced right-regulated system; for $\epsilon=\epsilon_{*}$ there is a unique proper ML-FP the $(\epsilon,0)$-reduced free system; for $\epsilon<\epsilon_{*}$ there is a unique ML-FP with proper left tail, but not proper on the right, for the $(\epsilon,0)$-reduced free system. $\Box$
@@ 9 Discussion
In this paper we do identify one case when “all possible nice properties hold,” namely the case when Condition 6.5 holds for any speed $v$. (In turn, Condition 6.5 necessarily holds when all job classes are $D$-monotone. And, in turn, if all classes have i.i.d. component sizes with IHR, this is sufficient for $D$-monotonicity.) Our results show that in this “nice” case: (a) $v_{min}=v_{max}=v_{*}$ and, therefore, the free ML-FP is unique and has speed $v_{*}$; (b) SSAI for the left-regulated system holds for the full range of loads; (c) the average advance velocity $v_{n}$ of the free system (with speed parameter 0) converges to $v_{*}$. Our results do not rule out that Condition 6.5 for any speed $v$ always holds for our general model, and then properties (a)-(c) always hold. Or, it is possible that Condition 6.5 does not always hold, but the nice properties (a)-(c) still hold always. Addressing these questions may be a subject of future work.